OpenStudy (anonymous):

Solve the quadratic equation 5x2 - 125 = 0

OpenStudy (anonymous):

would the two be an exponent?

OpenStudy (anonymous):

(oh and i love your icon. love that anime)

OpenStudy (anonymous):

x=+/- 5

OpenStudy (anonymous):

Right that anime was simply amazing, i found my awnser though c: thank you

OpenStudy (anonymous):

Actually the problem shuld be like this 5x^2 - 125 = 0 5x^2 - 125 = 0 5x^2 =125 x^2 =25 x= pm 5 (taking square root from both sides) pm= plus minus